# Matter-Mediated Entanglement in Classical Gravity: Suppression by Binding Potentials and Localization

### v1

@@ 1 Introduction
Aziz and Howl (hereafter AH) challenge the universality of this inference by arguing that, when matter is treated as a scalar field in a perturbative QFT framework, a classical gravitational field can still induce entanglement through higher-order processes involving virtual matter exchanging between the two systems [aziz2025classical]. On the other hand, several recent comments have questioned whether the specific nonrelativistic limit considered in AH actually generates entanglement, and emphasized that any entangling capability—if present—would originate from a quantum matter channel rather than from a purely classical mediator [marletto2025classicalgravitymediateentanglement, diósi2025noclassicalgravitydoes, sienicki2025commentclassicalgravityquantummatterclaimsgravitymediated].
@@ 2 QFT is not necessary
In particular, the diagram responsible for the classical-gravity-induced entanglement in AH (the upper-right diagram in Fig. 2 of [aziz2025classical], reproduced schematically in our Fig. 1) involves no particle creation or annihilation. Its physical content is a coherent matter-mediated amplitude connecting two spatial regions, which in quantum mechanics is precisely an evanescent/tunneling process. Put differently, what is called a “virtual particle” in a diagrammatic expansion corresponds, in a first-quantized description, to propagation through off-shell intermediate states (tunneling) rather than to an essentially field-theoretic effect.
@@ 3 Binding implies exponential suppression
In standard settings, tunneling/evanescent propagation over large distances is exponentially suppressed. This suppression is absent in AH’s process because binding/localization is not included in the microscopic model. Realistic matters are condensed many-body systems: microscopic constituents are localized by binding and lattice potentials with a large internal energy scale. In such systems, coherent propagation between separated bodies is evanescent and short-ranged.
@@
The quantum field–theoretic (QFT) description is consistent with this picture. A finite binding energy $E_{b}$, which raises the potential outside the masses, modifies the particle propagator across the bodies. This escaping energy can be interpreted as an effective mass shift in the propagator denominator,
@@
and we will consider it within the non-relativistic regime $E_{b}\ll mc^{2}$.
Because the mass shell itself is not shifted by $E_{b}$, this modification produces a mismatch in the exponential factors appearing in the relevant Feynman diagrams, which is central to AH’s analysis of distance-dependent decay. In their treatment, exponential factors arising from the contraction of the initial and final states cancel exactly against those of the propagator. Here, this cancellation fails, leaving a residual factor $e^{-\sqrt{m^{\prime 2}-m^{2}}\,\frac{c}{\hbar}d}\approx e^{-\sqrt{2mE_{b}}\,\frac{d}{\hbar}}$, which reproduces the suppression identified in the quantum-mechanical regime.
One can compare this result to Eq. (69) of AH’s Supplementary Material of [aziz2025classical], where our pole is fixed at $k_{0}=m^{\prime}c/\hbar$ rather than $mc/\hbar$. The absence of this decay in AH’s Eq. (69) arises from neglecting the role of the binding force in suppressing long-range propagation, treating it solely as a constraint on the initial and final states (Eq. (9) in the Supplementary Material). Incorporating a realistic $E_{b}$ restores an evanescent factor and renders the effective matter correlator short-ranged, scaling as $\sim e^{-d/\ell}$. Consequently, any matter-mediated entangling contribution requiring coherent propagation between macroscopic solids is exponentially suppressed in realistic platforms.
@@ 4 Quantum tunneling is not LOCC
This also clarifies why entanglement arises in AH’s analysis. The Feynman diagram Fig. 1 means that two objects can exchange particles. When particles are described by quantum mechanics, exchanging particles is beyond classical communication. We give a simple example in their two-object system. Suppose the first object has two energy levels $1L,1R$, and the second object has two energy levels $2L,2R$. Let the initial state be $|N_{1L}=N,N_{1R}=0\rangle\otimes|N_{2L}=N,N_{2R}=0\rangle$, and after a certain period of time, there is a small tunneling amplitude $\alpha\ll 1$ that an atom from $1L$ tunneling to $2R$. Then, up to normalization, the state becomes
$|\Psi\rangle\propto|N,0\rangle\otimes|N,0\rangle+\alpha\,|N-1,0\rangle\otimes|N,1\rangle,$ (4)
which is clearly entangled. This channel is essentially the same as AH’s virtual-matter channel, while it is simpler because no interaction is assumed. Even if there is no gravity, tunneling itself is able to generate entanglement. Nonetheless, this effect is exponentially suppressed in real experiments as long as the objects are solid or bounded.

### v2

@@ 1 Introduction
Aziz and Howl (hereafter AH) challenge the universality of this inference by arguing that, when matter is treated as a scalar field in a perturbative quantum field theory (QFT) framework, a classical gravitational field can still induce entanglement through higher-order processes involving virtual matter exchanging between the two systems [aziz_classical_2025]. On the other hand, several recent comments have questioned whether the specific nonrelativistic limit considered in AH actually generates entanglement, and emphasized that any entangling capability—if present—would originate from a quantum matter channel rather than from a purely classical mediator [marletto2025classicalgravitymediateentanglement, diósi2025noclassicalgravitydoes, sienicki2025commentclassicalgravityquantummatterclaimsgravitymediated].
@@ 2 QFT is not necessary
In particular, the diagram responsible for the classical-gravity-induced entanglement in AH (the upper-right diagram in Fig. 2 of [aziz_classical_2025], reproduced schematically in our Fig. 1) involves no particle creation or annihilation. Its physical content is a coherent matter-mediated amplitude connecting two spatial regions, which in quantum mechanics is precisely an evanescent/tunneling process. Put differently, what is called a “virtual particle” in a diagrammatic expansion corresponds, in a first-quantized description, to propagation through off-shell intermediate states (tunneling) rather than to an essentially field-theoretic effect.
@@ 3 Binding implies exponential suppression
In standard settings, tunneling/evanescent propagation over large distances is exponentially suppressed. This suppression is absent in AH’s process because binding/localization is not included in the microscopic model. Realistic matters are condensed many-body systems: microscopic constituents are localized by binding and lattice potentials with a large internal energy scale. In such systems, coherent propagation between separated bodies is evanescent and short-ranged. Even for one particle, a potential well is still necessary, otherwise the wavefunction will spread everywhere in the space.
@@
The QFT description is consistent with this picture. A finite binding energy $E_{b}$, which raises the potential outside the masses, modifies the particle propagator across the bodies. This escaping energy can be interpreted as an effective mass shift in the propagator (i.e. the mass parameter of $\frac{e^{ip\cdot x}}{p^{2}-m^{2}+i\varepsilon}$), which, after non-relativistic approximation $E_{b}\ll mc^{2}$, is
@@
This is similar to the energy conservation law, but can also be derived from solving the equation of green functions.
Notably the contraction with the initial and final states is not shifted by $E_{b}$, since according to the analysis in [aziz_classical_2025], these states are constricted inside the objects, where the binding potential vanishes (We have chosen the interior of the objects to be the zero-point of the binding potential, where the free mass parameter $m$ is defined). Hence, this modification produces a mismatch in different exponential factors appearing in the relevant Feynman diagrams, which is central to AH’s analysis of long-range decay. In their treatment, exponential factors arising from the contraction of the initial and final states cancel exactly against those of the propagator. Here, this cancellation fails, leaving a residual factor $e^{-\sqrt{m^{\prime 2}-m^{2}}\,\frac{c}{\hbar}d}\approx e^{-\sqrt{2mE_{b}}\,\frac{d}{\hbar}}$, which reproduces the suppression identified in the quantum-mechanical regime.
One can compare this result to Eq. (69) of AH’s Supplementary Material of [aziz_classical_2025], where our pole is fixed at $k_{0}=m^{\prime}c/\hbar$ rather than $mc/\hbar$. The absence of this decay in AH’s Eq. (69) arises from neglecting the role of the binding force in suppressing long-range propagation, treating it solely as a constraint on the initial and final states (Eq. (9) in the Supplementary Material). Incorporating a realistic $E_{b}$ restores an evanescent factor and renders the effective matter correlator short-ranged, scaling as $\sim e^{-d/\ell}$. Consequently, any matter-mediated entangling contribution requiring coherent propagation between macroscopic solids is exponentially suppressed in realistic platforms.
@@ 4 Quantum tunneling is not LOCC
This also clarifies why entanglement arises in AH’s analysis. The Feynman diagram Fig. 1 means that two objects can exchange particles. When particles are described by quantum mechanics, exchanging particles is beyond classical communication. For example, we assume that in the lower-left object (bucket) of Fig. 1, the purple ball and a red ball are entangled. This entanglement is local, but after the quantum tunneling, the entanglement will be between the two macroscopic objects.
This channel is essentially the same as AH’s virtual-matter channel, while it is simpler because no interaction is assumed. Even if there is no gravity, tunneling itself is able to generate entanglement. This effect is exponentially suppressed in real experiments as long as the objects are solid or bounded.